1. Use the confidence interval to find the estimated margin of error. Then find the sample mean. A biologist reports a confidence interval of (4.0, 4.8) when estimating the mean height (in centimeters) of a sample of seedlings.

2. In a random sample of 26 people, the mean commute time to work was 32.2 minutes and the standard deviation was 7.1 minutes. Assume the population is normally distributed and use a t-distribution to construct a 80% confidence interval for the population mean "u". What is the margin of error of "u"? Interpret the results.
Answer #1

1)

margin of error = (upper bound - lower bound)/2 = (4.8 -4)/2 = 0.4

sample = (upper bound + lower bound)/2 = (4.8 +4)/2 = 4.4

Please rate

Please post next question again